# Classification of abelian Schur groups I

### v1

@@ 6.2. $S$-rings over cyclic $p$-groups
Let $L=\operatorname{rad}(X)$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/L$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muT\mkern-1.5mu}\mkern 1.5mu$. The $S$-ring $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ has a trivial radical. So $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ is trivial or cyclotomic by Lemma 6.2. As $X$ is not regular, $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$ also is not regular. Therefore $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ can not be cyclotomic and hence it is trivial. Thus, $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\cup\{L\}$ is an $\mathcal{A}_{G/L}$-subgroup. Since $L=\operatorname{rad}(X)$, we conclude that $X=U\setminus L$ for some $\mathcal{A}$-subgroup $U$ as required.
Now let us prove the second part of the lemma. If $U=G$, then $\mathcal{A}$ is the $L/L$-wreath product (possibly, trivial) and the trivial $U/U$-wreath product. Further, we assume that $U<G$ and hence $\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu<\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu$. The latter implies that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial. Note that $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu=(\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu)^{\#}$ is a nonregular basic set of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ with trivial radical. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ can not be cyclotomic and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is a nontrivial generalized wreath product by Lemma 6.2. In view of Lemma 6.10, $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is the nontrivial $W/V$-wreath product, where $\operatorname{rad}(\mathcal{A}_{W})$ is trivial and $V$ is the least nontrivial $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$-subgroup. Since $\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)$ is trivial, we conclude that $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\subseteq W$. This yields that $\mathcal{A}_{W}$ is not cyclotomic. Due to Lemma 6.2, the $S$-ring $\mathcal{A}_{W}$ is trivial. Therefore $W=V=\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu$. Thus,
@@
In view of the above equality, to prove the lemma it is enough to show that $L\leq\operatorname{rad}(Y)$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. Assume the contrary that $N=\operatorname{rad}(Y)<L$ for some $Y\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/N$ is denoted by $\widetilde{T}$. As $Y$ lies outside $U$, we conclude that $X\subseteq U\leq\langle Y\rangle$. So $\widetilde{X}$ is a nonregular basic set of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$. This implies that $\mathcal{A}_{\langle\widetilde{Y}\rangle}$ can not be cyclotomic. Therefore $\mathcal{A}_{\langle\widetilde{Y}\rangle}$ is trivial by Lemma 6.2. On the other hand, $\widetilde{X}$ and $\widetilde{Y}$ are nontrivial basic sets of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$, a contradiction to triviality of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$. ∎
@@ 6.3. $S$-rings over some other cyclic groups
An $S$-ring over a group of prime order is cyclotomic by Lemma 6.7 and hence a tensor product of $S$-rings over groups of prime orders is cyclotomic by Eq. (4). Therefore the required follows from Lemma 6.2. ∎
@@
Assume that $\mathcal{A}$ is nonnormal. Then $\mathcal{A}=\mathbb{Z}_{L}\otimes\mathcal{A}_{U}$ by Lemma 6.2. If $|\operatorname{rad}(\mathcal{A}_{U})|>1$, then $|\operatorname{rad}(\mathcal{A})|>1$, a contradiction to the assumption of the lemma. So $|\operatorname{rad}(\mathcal{A}_{U})|=1$. If $\mathcal{A}_{U}\neq\mathcal{T}_{U}$, then $\mathcal{A}_{U}$ is normal by Lemma 6.2 and consequently $\mathcal{A}$ so is by Eq. (3), a contradiction to the assumption of the lemma. Thus, $\mathcal{A}_{U}=\mathcal{T}_{U}$ and hence $\mathcal{A}=\mathbb{Z}_{L}\otimes\mathcal{T}_{U}$. ∎
@@ 7. $S$-rings over $E_{4}\times C_{n}$
    $\mathcal{A}$ is dense, $U\geq H$, $\mathcal{A}_{U}$ is cyclotomic, $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=1$ unless $n=3^{k}$ and $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=3$, and $L\leq D$.
Observe that $U\cap D$ in Statement $(4)$ of Theorem 7.1 is cyclic and hence $\operatorname{rad}(\mathcal{A}_{U\cap D})$ is defined correctly. Before the proof of Theorem 7.1, we need some preparatory work. In the first subsection, we study a structure of a basic set of $\mathcal{A}$. The second and third subsections deal with the cases when $\mathcal{A}$ is nondense and when $\mathcal{A}$ is dense, respectively. In the fourth subsection, we prove Theorem 7.1. The fifth subsection contains some auxiliary statements which will be used in the proof of Theorem 1.2.
Throughout this section, $p$ and $q$ are odd primes, $k$ is a positive integer, $n\in\{p^{k},pq\}$, $D\cong C_{n}$, $P=D_{p}=\{g\in D:~|g|\in\{1,p\}\}$, $H\cong E_{4}$, $G=H\times D$, and $\mathcal{A}$ is an $S$-ring over $G$.
@@ 7.1. Structure of a basic set
Assume that $P\nleq\operatorname{rad}(X)$, i.e. Statement $(1)$ of the lemma does not hold. Then the union $X^{0}$ of all $hX_{h,l}$, $h\in H$, $l\in[n]$, such that $X_{h,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{h,l})$, is nonempty. As $D\cong C_{p^{k}}$, the condition $P\nleq\operatorname{rad}(X_{h,l})$ is equivalent to $|\operatorname{rad}(X_{h,l})|=1$. Lemma 2.5 implies that each $X_{h,l}$ is an orbit of some subgroup $K$ of $\operatorname{Aut}(D_{l})$. From Lemma 6.8(1) it follows that $|K|\leq p-1$ and hence
@@
for every $h\in H$ and $l\in[n]$ with $X_{h,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{h,l})$. Put $H_{0}=(X^{0})_{H}$ and $X^{i}=(X^{i-1})^{[p]}$ for $i\geq 1$. Due to the definition of $X^{0}$ and Eq. (5), we have $X^{1}=X^{[p]}$ and $X^{i}=\{x^{p}:~x\in X^{i-1}\}$. Therefore $(X^{i})_{H}=H_{0}$ for every $i\geq 1$, in particular, $(X^{1})_{H}=(X^{[p]})_{H}=H_{0}$. By Lemma 2.4, the set $X^{i}$ is an $\mathcal{A}$-set for every $i\geq 1$. So $X^{k}=H_{0}$ is an $\mathcal{A}$-set. Thus, $\langle H_{0}\rangle$ is an $\mathcal{A}$-subgroup. If $H_{0}\neq\{e\}$, then Statement $(2)$ of the lemma holds.
@@
$X^{0}\subseteq X\cap D\neq\varnothing.$
Let $l=p^{j}$ be the greatest integer such that $X_{e,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{e,l})$. By the assumption of the lemma, $X\setminus D\neq\varnothing$. So
@@
If $j\neq 1$, then $X^{j-1}$ is a nonemepty $\mathcal{A}$-set inside $P^{\#}$. Therefore $P=\langle X^{j-1}\rangle$ is an $\mathcal{A}$-subgroup. Now Lemma 2.1 and Eq. (6) imply that $P\leq\operatorname{rad}(X)$, a contradiction to the assumption. If $j=1$, then $X_{e,l}\subseteq P^{\#}$. Eq. (6) and Lemma 2.2 applied to $X$ and $P$ yield that $\operatorname{rad}(X)\leq P$ and $X=\langle X\rangle\setminus\operatorname{rad}(X)$. By the assumption, $P\nleq\operatorname{rad}(X)$ and consequently $\operatorname{rad}(X)$ is trivial. Thus, $X=\langle X\rangle\setminus\{e\}$ and Statement $(3)$ of the lemma holds. ∎
@@ 7.2. Nondense $S$-rings
Proposition 7.3.
Suppose that $H$ is not an $\mathcal{A}$-subgroup. Then $\mathcal{A}$ is trivial, or a nontrivial tensor product, or a nontrivial $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ satisfying one of the following conditions:
@@
    $4\nmid|U|$ and $|L|\neq 2$.
@@
Let $n=pq$. Observe that the subgroup of $G$ of order $4p$ or the subgroup of $G$ of order $4q$ is not an $\mathcal{A}$-subgroup because otherwise $H$ is an $\mathcal{A}$-subgroup as an intersection of these subgroups which contradicts to the assumption of the proposition. Without loss of generality, we may assume that the subgroup of $G$ of order $4q$ is not an $\mathcal{A}$-subgroup. Then
@@
Let $P_{1}$ be the least $\mathcal{A}$-subgroup containing $P\cong C_{p}$. One of the statements of Lemma 5.3 holds for $\mathcal{A}$, $H_{1}$, and $P_{1}$. If Statement $(1)$ of Lemma 5.3 holds, then $H_{1}P_{1}=G$ and $\mathcal{A}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$. If $H_{1}\cap P_{1}$ is trivial, then $\mathcal{A}=\mathcal{A}_{H_{1}}\otimes\mathcal{A}_{P_{1}}$ is a nontrivial tensor product and we are done. Otherwise $\mathcal{A}$ is the nontrivial $U/L=H_{1}/(H_{1}\cap P_{1})$-wreath product. Together with Eq. (7), this implies that $U/L$ satisfies Condition $(1)$ from Proposition 7.3 as desired or $|H_{1}|=2q$ and $|H_{1}\cap P_{1}|=2$. In the latter case, Lemma 6.12 yields that $\mathcal{A}_{H_{1}}=\mathcal{A}_{H_{1}\cap P_{1}}\otimes\mathcal{A}_{Q}$, where $Q$ is an $\mathcal{A}$-subgroup of order $q$, or $\mathcal{A}_{H_{1}}=\mathcal{A}_{H_{1}\cap P_{1}}\wr\mathcal{A}_{H_{1}/(H_{1}\cap P_{1})}$. In the first case, $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ and $U/L$ satisfies Condition $(2)$. In the second one, $\mathcal{A}$ is the $L/L=(H_{1}\cap P_{1})/(H_{1}\cap P_{1})$-wreath product and the section $L/L$ satisfies Condition $(1)$ .
@@
If Statement $(2)$ of Lemma 5.3 holds, then $\mathcal{A}$ is trivial whenever $H_{1}$ is trivial and the $U/L=H_{1}/H_{1}$-wreath product satisfying Condition $(1)$ otherwise. If Statement $(3)$ of Lemma 5.3 holds, then $\mathcal{A}$ is the nontrivial $U/L=(H_{1}P_{1})/P_{1}$-wreath product. Due to Eq. (7), the section $U/L$ satisfies Condition $(1)$ or $(3)$.
@@
Now let $n=p^{k}$. As in the previous paragraph, $P_{1}$ denotes the least $\mathcal{A}$-subgroup containing $P$. One of Statements (1)(1)-$(3)$ of Lemma 7.2 holds for every basic set of $\mathcal{A}$ outside $D$. We divide the rest of the proof into two cases depending on existence of a basic set outside $D$ satisfying Statement $(3)$ of Lemma 7.2.
@@
Case 1. Suppose that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus D}$ satisfying Statement $(3)$ of Lemma 7.2, i.e. $X\cup\{e\}$ is an $\mathcal{A}$-subgroup with $X\cap D\neq\varnothing$. Then $X\cup\{e\}=P_{1}$, $P_{1}\nleq D$, and $\mathcal{A}_{P_{1}}=\mathcal{T}_{P_{1}}$. If $P_{1}=G$, then $\mathcal{A}$ is trivial. Further, we assume that $G\setminus P_{1}\neq\varnothing$.
@@
Observe that $Y\nsubseteq D$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$. Indeed, otherwise $\langle Y\rangle$ is an $\mathcal{A}$-subgroup inside $D$. So $\langle Y\rangle\geq P_{1}$ and hence
$D\geq\langle Y\rangle\geq P_{1}\nleq D,$
a contradiction. Thus, Lemma 7.2 holds for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$.
@@
Statement $(3)$ of Lemma 7.2 can not hold for $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$ because $P_{1}$ is the least $\mathcal{A}$-subgroup containing $P$. If Statement $(1)$ of Lemma 7.2 holds for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$, then $P_{1}\leq\operatorname{rad}(Y)$ and hence $\mathcal{A}$ is the $U/L=P_{1}/P_{1}$-wreath product and $U/L$ satisfies Condition $(1)$ from Proposition 7.3.
@@
In view of the above paragraph, we may assume that there is $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$ satisfying Statement $(2)$ of Lemma 7.2 and $P_{1}\nleq\operatorname{rad}(Y)$. Since $H$ is not an $\mathcal{A}$-subgroup, $A=\langle(Y^{[p]})_{H}\rangle$ is an $\mathcal{A}$-subgroup of order $2$. Clearly, $\mathcal{A}_{A}=\mathbb{Z}A$. Note that $A\cap P_{1}$ is trivial because $\mathcal{A}_{P_{1}}=\mathcal{T}_{P_{1}}$. Therefore
(8) $\mathcal{A}_{A\times P_{1}}=\mathbb{Z}A\otimes\mathcal{T}_{P_{1}}.$

### v3

@@ 6.2. $S$-rings over cyclic $p$-groups
Let $L=\operatorname{rad}(X)$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/L$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muT\mkern-1.5mu}\mkern 1.5mu$. The $S$-ring $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ has a trivial radical. So $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ is trivial or cyclotomic by Lemma 6.2. As $X$ is not regular, $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$ also is not regular. Therefore $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ can not be cyclotomic and hence it is trivial. Thus, $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\cup\{L\}$ is an $\mathcal{A}_{G/L}$-subgroup. Since $L=\operatorname{rad}(X)$, we conclude that $X=U\setminus L$ for some $\mathcal{A}$-subgroup $U$ as required. One can see that $|U:L|\geq p^{2}$ because otherwise $X$ is regular which contradicts to the assumption of the lemma.
Now let us prove the second part of the lemma. If $U=G$, then $\mathcal{A}$ is the $L/L$-wreath product (possibly, trivial) and the trivial $U/U$-wreath product. Further, we assume that $U<G$ and hence $\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu<\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu$. The latter implies that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial. Note that $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu=(\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu)^{\#}$ is a nonregular basic set of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ with trivial radical. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ can not be cyclotomic and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is a nontrivial generalized wreath product by Lemma 6.2. In view of Lemma 6.10, $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is the nontrivial $W/V$-wreath product, where $\operatorname{rad}(\mathcal{A}_{W})$ is trivial and $V$ is the least nontrivial $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$-subgroup. Since $\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)$ is trivial, we conclude that $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\subseteq W$. Hence $\mathcal{A}_{W}$ is not cyclotomic. Due to Lemma 6.2, the $S$-ring $\mathcal{A}_{W}$ is trivial. Therefore $W=V=\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu$. Thus,
@@
In view of the above equality, to prove the lemma it suffices to show that $L\leq\operatorname{rad}(Y)$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. Assume the contrary that $N=\operatorname{rad}(Y)<L$ for some $Y\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/N$ is denoted by $\widetilde{T}$. As $Y$ lies outside $U$, we have $X\subseteq U\leq\langle Y\rangle$. So $\widetilde{X}$ is a nonregular basic set of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$. This implies that $\mathcal{A}_{\langle\widetilde{Y}\rangle}$ can not be cyclotomic. Therefore $\mathcal{A}_{\langle\widetilde{Y}\rangle}$ is trivial by Lemma 6.2. On the other hand, $\widetilde{X}$ and $\widetilde{Y}$ are nontrivial basic sets of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$, a contradiction. ∎
@@ 6.3. $S$-rings over some other cyclic groups
An $S$-ring over a group of prime order is cyclotomic by Lemma 6.7 and hence a tensor product of $S$-rings over groups of prime orders is cyclotomic by Eq. (4). Therefore the required statement follows from Lemma 6.2. ∎
@@
Assume that $\mathcal{A}$ is nonnormal. Then $\mathcal{A}=\mathbb{Z}_{L}\otimes\mathcal{A}_{U}$ by Lemma 6.2. If $|\operatorname{rad}(\mathcal{A}_{U})|>1$, then $|\operatorname{rad}(\mathcal{A})|>1$, a contradiction to the assumption of the lemma. So $|\operatorname{rad}(\mathcal{A}_{U})|=1$. If $\mathcal{A}_{U}\neq\mathcal{T}_{U}$, then $\mathcal{A}_{U}$ is normal by Lemma 6.2 and consequently so is $\mathcal{A}$ by Eq. (3), a contradiction to the assumption of the lemma. Thus, $\mathcal{A}_{U}=\mathcal{T}_{U}$ and hence $\mathcal{A}=\mathbb{Z}_{L}\otimes\mathcal{T}_{U}$. ∎
@@ 7. $S$-rings over $E_{4}\times C_{n}$
    $\mathcal{A}$ is dense, $L\leq D$, $U\geq H$, $\mathcal{A}_{U}$ is cyclotomic, and $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=1$ unless $n=3^{k}$ and $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=3$.
Observe that $U\cap D$ in Statement $(4)$ of Theorem 7.1 is cyclic and hence $\operatorname{rad}(\mathcal{A}_{U\cap D})$ is defined correctly.
Clearly, Theorem 1.3 immediately follows from Theorem 7.1.
We divide the proof of Theorem 7.1 into two cases depending on whether $\mathcal{A}$ is nondense or dense. We prove that every nondense nontrivial $S$-ring over $G$ is a nontrivial tensor or generalized wreath product and every dense $S$-ring over $G$ is cyclotomic or a nontrivial generalized wreath product. The nondense case is divided into two subcases depending on whether $H$ or $D$ is not an $\mathcal{A}$-subgroup (Theorems 7.3 and 7.8, respectively), whereas the dense case is divided into two subcases depending on whether $\operatorname{rad}(\mathcal{A}_{D})$ is trivial or not (Theorems 7.12 and 7.13, respectively).
In the first subsection, we study a structure of a basic set of $\mathcal{A}$. The second and third subsections deal with the above mentioned cases when $\mathcal{A}$ is nondense and when $\mathcal{A}$ is dense, respectively. In the fourth subsection, we give a proof of Theorem 7.1. The fifth subsection contains some auxiliary statements to be used in the proof of Theorem 1.2.
Throughout this section, $p$ and $q$ are odd primes, $k$ is a positive integer, $n\in\{p^{k},pq\}$, $D\cong C_{n}$, $P=D_{p}=\{g\in D:~g^{p}=e\}$, $H\cong E_{4}$, $G=H\times D$, and $\mathcal{A}$ is an $S$-ring over $G$.
@@ 7.1. Structure of a basic set
Assume that $P\nleq\operatorname{rad}(X)$, i.e. Statement $(1)$ of the lemma does not hold. Then the union $Y$ of all $hX_{h,l}=X\cap hD_{l}^{*}$, $h\in H$, $l\in\operatorname{Div}(n)$, such that $X_{h,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{h,l})$ is nonempty. As $D\cong C_{p^{k}}$ and $X_{h,l}\subseteq D$, the condition $P\nleq\operatorname{rad}(X_{h,l})$ is equivalent to $|\operatorname{rad}(X_{h,l})|=1$. Lemma 2.5 implies that each $X_{h,l}$ is an orbit of some subgroup $K$ of $\operatorname{Aut}(D_{l})$. From Lemma 6.8(1) it follows that $|K|\leq p-1$ and hence
@@
for all $h\in H$ and $l\in\operatorname{Div}(n)$ such that $X_{h,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{h,l})$.
Put $H_{0}=Y_{H}$, $Y^{[p]^{0}}=Y$, $Y^{[p]^{1}}=Y^{[p]}$, and $Y^{[p]^{i}}=(Y^{[p]^{i-1}})^{[p]}$ for $i\geq 2$. Due to the definition of $Y$ and Eq. (5), we have $Y^{[p]}=X^{[p]}$ and $Y^{[p]^{i}}=\{y^{p}:~y\in Y^{[p]^{i-1}}\}$. Therefore $(Y^{[p]^{i}})_{H}=H_{0}$ for every $i\geq 1$, in particular, $(Y^{[p]})_{H}=(X^{[p]})_{H}=H_{0}$. By Lemma 2.4, the set $Y^{[p]^{i}}$ is an $\mathcal{A}$-set for every $i\geq 1$. So $Y^{[p]^{k}}=H_{0}$ is an $\mathcal{A}$-set. Thus, $\langle H_{0}\rangle$ is an $\mathcal{A}$-subgroup. If $H_{0}\neq\{e\}$, then Statement $(2)$ of the lemma holds.
@@
$Y\subseteq X\cap D\neq\varnothing.$
Let $l=p^{j}$ be the greatest integer such that $X_{e,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{e,l})$. By the supposition of the lemma, $X\setminus D\neq\varnothing$. So
@@
If $j\neq 1$, then $Y^{[p]^{j-1}}$ is a nonemepty $\mathcal{A}$-set inside $P^{\#}$. Therefore $P=\langle Y^{[p]^{j-1}}\rangle$ is an $\mathcal{A}$-subgroup. Now Lemma 2.1 and Eq. (6) imply that $P\leq\operatorname{rad}(X)$ in contrast to the assumption $P\nleq\operatorname{rad}(X)$. If $j=1$, then $X_{e,l}\subseteq P^{\#}$. Eq. (6) and Lemma 2.2 applied to $X$ and $P$ yield that $\operatorname{rad}(X)\leq P$ and $X=\langle X\rangle\setminus\operatorname{rad}(X)$. By the assumption, $P\nleq\operatorname{rad}(X)$ and consequently $\operatorname{rad}(X)$ is trivial. Thus, $X=\langle X\rangle\setminus\{e\}$ and Statement $(3)$ of the lemma holds. ∎
@@ 7.2. Nondense $S$-rings
Theorem 7.3.
Suppose that $H\cong E_{4}$ is not an $\mathcal{A}$-subgroup. Then $\mathcal{A}$ is trivial, or a nontrivial tensor product, or a nontrivial $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ satisfying one of the following conditions:
@@
    $4\nmid|U|$ and $|\operatorname{rad}(\mathcal{A}_{U})|=1$.
We divide the proof of Theorem 7.3 into two propositions.
Proposition 7.4.
If $H$ is not an $\mathcal{A}$-subgroup, then Theorem 7.3 holds unless $\mathcal{A}$ is a nontrivial $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ such that $4\nmid|U|$ and $|L|\neq 2$.
@@
Let $n=pq$. Observe that the subgroup of $G$ of order $4p$ or the subgroup of $G$ of order $4q$ is not an $\mathcal{A}$-subgroup because otherwise $H$ is an intersection of $\mathcal{A}$-subgroups which contradicts to the supposition of the proposition. Without loss of generality, we may assume that the subgroup of $G$ of order $4q$ is not an $\mathcal{A}$-subgroup. Then
@@
Let $P_{1}$ be the least $\mathcal{A}$-subgroup containing $P\cong C_{p}$. One of the statements of Lemma 5.3 holds for $\mathcal{A}$, $H_{1}$, and $P_{1}$. If Statement $(1)$ of Lemma 5.3 holds, then $H_{1}P_{1}=G$ and $\mathcal{A}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$. If $H_{1}\cap P_{1}$ is trivial, then $\mathcal{A}=\mathcal{A}_{H_{1}}\otimes\mathcal{A}_{P_{1}}$ is a nontrivial tensor product and we are done. Otherwise, $\mathcal{A}$ is the nontrivial $U/L=H_{1}/(H_{1}\cap P_{1})$-wreath product. Together with Eq. (7), this implies that $|U/L|\leq 2$ as desired or $|H_{1}|=2q$ and $|H_{1}\cap P_{1}|=2$. In the latter case, Lemma 6.12 yields that $\mathcal{A}_{H_{1}}=\mathcal{A}_{H_{1}\cap P_{1}}\otimes\mathcal{A}_{Q}$, where $Q$ is an $\mathcal{A}$-subgroup of order $q$, or $\mathcal{A}_{H_{1}}=\mathcal{A}_{H_{1}\cap P_{1}}\wr\mathcal{A}_{H_{1}/(H_{1}\cap P_{1})}$. In the first case, $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ and $U/L$ satisfies Condition $(2)$ from Theorem 7.3 as required. In the second one, $\mathcal{A}$ is the $L/L=(H_{1}\cap P_{1})/(H_{1}\cap P_{1})$-wreath product and the section $L/L$ satisfies Condition $(1)$ from Theorem 7.3 as required.
@@
If Statement $(2)$ of Lemma 5.3 holds, then $\mathcal{A}$ is trivial as desired whenever $H_{1}$ is trivial and the $U/L=H_{1}/H_{1}$-wreath product satisfying Condition $(1)$ from Theorem 7.3 otherwise. If Statement $(3)$ of Lemma 5.3 holds, then $\mathcal{A}$ is the nontrivial $U/L=(H_{1}P_{1})/P_{1}$-wreath product. Due to Eq. (7), the section $U/L$ satisfies Condition $(1)$ from Theorem 7.3 or $4\nmid|U|$ and $|L|\neq 2$ as required.
@@
Now let $n=p^{k}$, $k\geq 1$. As in the previous paragraph, $P_{1}$ denotes the least $\mathcal{A}$-subgroup containing $P$. One of Statements (1)(1)-$(3)$ of Lemma 7.2 holds for every basic set of $\mathcal{A}$ outside $D$. We divide the rest of the proof into two cases depending on existence of a basic set outside $D$ satisfying Statement $(3)$ of Lemma 7.2.
@@
Case 1. Suppose that there is a basic set of $\mathcal{A}$ outside $D$, satisfying Statement $(3)$ of Lemma 7.2. Then $\mathcal{A}_{P_{1}}=\mathcal{T}_{P_{1}}$ and $P_{1}\nleq D$. If $P_{1}=G$, then $\mathcal{A}$ is trivial and we are done. Further, we assume that $G\setminus P_{1}\neq\varnothing$.
@@
Observe that the condition of Lemma 7.2 holds for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$, i.e. $X\nsubseteq D$. Indeed, otherwise $\langle X\rangle$ is an $\mathcal{A}$-subgroup inside $D$. So $\langle X\rangle\geq P_{1}$ and hence
$D\geq\langle X\rangle\geq P_{1}\nleq D,$
a contradiction.
@@
The above paragraph implies that one of Statements (1)(1)-$(3)$ from Lemma 7.2 holds for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$. Statement $(3)$ can not hold for $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$ because $P_{1}$ is a unique minimal $\mathcal{A}$-subgroup containing $P$. If Statement $(1)$ holds for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$, then $P_{1}\leq\operatorname{rad}(X)$ and hence $\mathcal{A}$ is the $U/L=P_{1}/P_{1}$-wreath product and $U/L$ satisfies Condition $(1)$ from Theorem 7.3 as required.
@@
In view of the previous paragraph, we may assume that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$ for which Statement $(2)$ of Lemma 7.2 holds and $P_{1}\nleq\operatorname{rad}(X)$. Since $H$ is not an $\mathcal{A}$-subgroup, $A=\langle(X^{[p]})_{H}\rangle$ is an $\mathcal{A}$-subgroup of order $2$. Clearly, $\mathcal{A}_{A}=\mathbb{Z}A$. Note that $A\cap P_{1}$ is trivial because $\mathcal{A}_{P_{1}}=\mathcal{T}_{P_{1}}$. Therefore
(8) $\mathcal{A}_{A\times P_{1}}=\mathbb{Z}A\otimes\mathcal{T}_{P_{1}}$